in a compound of magnesium and nitrogen 54g of magnesium combined with21g of nitrogen.determine the emphirical formulae of te compound



1
Expert's answer
2021-05-04T14:16:27-0400

Moles of Magnesium "= \dfrac{54}{24} = 2.25"


Moles of Nitrogen "= \dfrac{21}{14} = 1.5"


Relative number of moles of atom,


For Magnesium "= \dfrac{2.25}{1.5} = 1.5"


For Nitrogen "= \dfrac{1.5}{1.5} = 1"


Converting numbers into whole numbers.


Hence, Empirical Formula will be "= Mg_{3}N_2"